# Towards a Common Framework for Autoformalization

### v1

@@ 1 Introduction
Drawing on the long tradition of formalizing human reasoning through mathematics and logic, AI pioneers (???) incorporated ideas from formal systems into functional and logical programming languages for formalizing knowledge and solving practical problems. Building on these foundations, researchers in natural language processing (?) and language understanding (?) began applying formal representations and inference techniques to human language (?). This enabled computers to reason about the meaning of sentences, answer questions, and interact with users in more human-like ways, thus bridging the gap between symbolic AI and language understanding. These efforts opened the way in the development of semantic parsing (?), a field dedicated to translating natural language into formal meaning representations.
The term Autoformalization was originally introduced to describe the automatic formalization of mathematics with interactive theorem provers (proof assistants) such as Isabelle/HOL, Lean, Rocq, and Mizar, a field that has grown rapidly in recent years; there is increasing interest in employing deep learning to automate the task of formalization (????). Earlier, pioneering work employing techniques from machine learning had introduced the automation of the task of formalization referring to it as “corpus-based translation between informal and formal mathematics” (?) and “parsing between aligned corpora”, exploring the challenge: Is it possible to automatically parse informal mathematical texts into formal ones and formally verify them? (?). Autoformalization is closely related to semantic parsing (?) and can be seen as an instance of semantic parsing. The distinctive feature of autoformalization would be that—unlike semantic parsing which encompasses a broad range of target formalisms including general purpose programming languages—it is restricted to target formal languages used in logical inference and automated reasoning. Independently of this distinction, with the development of deep learning approaches to natural language processing—particularly the emergence of transformer architectures—the task of translating informal language into formal representations is most often performed using large language models (LLMs) [^1] The commonly used models include Claude, Codex, DeepSeek, GPT3.5, GPT4, Llemma, Minerva, Mistral, and Mixtral.. This approach is more general and flexible, and less tied to a specific target language or domain than traditional semantic parsing methods; however, it comes at the cost of introducing new challenges in verifying their less restricted outputs.
@@ 2 Review of Existing Usage and Definitions
The term autoformalization originated in the context of formalizing mathematics in 2018 (?), and as a result, it is widely adopted within this domain. Since the recent emergence of LLMs, users of proof assistants have been experimenting with using LLMs to assist with the formalization of mathematics, yielding mixed but potentially promising results (?). All 32 papers reviewed for a definition from this subfield use the term autoformalization. Most of works provide only informal definitions of autoformalization, which may reflect the fact that formalization of mathematics is intuitively well understood within the community and does not require defining. These informal definitions typically describe autoformalization as the task of automatically translating mathematical content from informal language into formal languages (??????????????????). Some authors also explicitly include machine-verifiability of the output as part of the definition (???????????). Only one group of authors attempted a formal definition (??).
@@
In this category, we reviewed 30 papers. The target formalisms include first order logic (?????????????); declarative programming language, Prolog (???) and answer set programming (ASP) (????). Another main line of work focuses on translating problems from natural language into temporal logics such as LTL (?) and CTL (?). These languages extend classical logic with temporal operators to reason about properties that hold over time, with LTL describing linear sequences of states and CTL describing branching time structures where multiple futures are possible from each state. There exist a significant number of recent works focused on translating natural language descriptions to these languages (??????????); one the reasons of popularity of temporal logics is its relevance in industrial applications.
Only 8 of reviewed papers explicitly used the term autoformalization, and all of them defined it informally—typically as the translation of informal or natural language into formal representations (????????). Among the remaining works, one paper formally defines the task (?), while the rest describe it informally as translation from natural language into a specific formalism. Notably, even though these papers do not use the term autoformalization, their definitions are consistent with those that do.
@@
In the planning category, we included 16 papers. None of them use the term “autoformalization”. One of the works provides a formal definition of the task (?) as defining a mapping function generating a world model from a natural language description. The rest of the works define the task informally (??????????????). In some cases the task may encompass generating both domain and problem files (?), while in others only PDDL snippets are being instantiated, e.g., goals in  (?). A related but separate line of research involves using LLMs as planners themselves, generating plans from PDDL inputs. This direction falls outside the scope of the present work.
@@
Several works that we surveyed (????) focus on translating natural language descriptions of real-world domains into ontologies using the Web Ontology Language (OWL) (?), a formal language used to define and share structured, machine-readable knowledge in terms of individuals, concepts, and relationships between them. Some other works produce ontologies expressed in RDF or RDFS (?). RDF allows us to express Knowledge Graphs, where nodes represent named entities, node labels represent their individual properties (called “types” or “classes”), and labelled edges represent relations between them. RDFS enriches RDF with information about subsumption, domain, and range relations between classes. Both RDF and RDFS can be seen as subsets of OWL. Works such as (??) translate natural language to RDFS and RDF ontologies, respectively. Others express ontologies in ad-hoc languages (??), which seem to be expressible in OWL or RDFS. In all these works, syntactic validity is checked automatically, and semantic is checked manually, or by running certain conjunctive queries on the produced ontology. None of the reviewed works in this category employs the term autoformalization to denote the translation task.
@@ 4 Challenges and Open Problems
While there is a growing body of work aimed at advancing autoformalization (?????????), several fundamental challenges still offer exciting research opportunities that will shape the field’s future development. We next provide a selective overview of these challenges and opportunities.

### v2

@@ 1 Introduction
Drawing on the long tradition of formalizing human reasoning through mathematics and logic, AI pioneers (?; ?; ?) incorporated ideas from formal systems into functional and logical programming languages for formalizing knowledge and solving practical problems. Building on these foundations, researchers in natural language processing (?) and language understanding (?) began applying formal representations and inference techniques to human language (?). This enabled computers to reason about the meaning of sentences, answer questions, and interact with users in more human-like ways, thus bridging the gap between symbolic AI and language understanding. These efforts opened the way in the development of semantic parsing (?), a field dedicated to translating natural language into formal meaning representations.
The term Autoformalization was originally introduced to describe the automatic formalization of mathematics with interactive theorem provers (proof assistants) such as Isabelle/HOL, Lean, Rocq, and Mizar, a field that has grown rapidly in recent years; there is increasing interest in employing deep learning to automate the task of formalization (?; ?; ?; ?). Earlier, pioneering work employing techniques from machine learning had introduced the automation of the task of formalization referring to it as “corpus-based translation between informal and formal mathematics” (?) and “parsing between aligned corpora”, exploring the challenge: Is it possible to automatically parse informal mathematical texts into formal ones and formally verify them? (?). Autoformalization is closely related to semantic parsing (?) and can be seen as an instance of semantic parsing. The distinctive feature of autoformalization would be that—unlike semantic parsing which encompasses a broad range of target formalisms including general purpose programming languages—it is restricted to target formal languages used in logical inference and automated reasoning. Independently of this distinction, with the development of deep learning approaches to natural language processing—particularly the emergence of transformer architectures—the task of translating informal language into formal representations is most often performed using large language models (LLMs) [^1] The commonly used models include Claude, Codex, DeepSeek, GPT3.5, GPT4, Llemma, Minerva, Mistral, and Mixtral.. This approach is more general and flexible, and less tied to a specific target language or domain than traditional semantic parsing methods; however, it comes at the cost of introducing new challenges in verifying their less restricted outputs.
@@ 2 Review of Existing Usage and Definitions
The term autoformalization originated in the context of formalizing mathematics in 2018 (?), and as a result, it is widely adopted within this domain. Since the recent emergence of LLMs, users of proof assistants have been experimenting with using LLMs to assist with the formalization of mathematics, yielding mixed but potentially promising results (?). All 32 papers reviewed for a definition from this subfield use the term autoformalization. Most of works provide only informal definitions of autoformalization, which may reflect the fact that formalization of mathematics is intuitively well understood within the community and does not require defining. These informal definitions typically describe autoformalization as the task of automatically translating mathematical content from informal language into formal languages (?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?). Some authors also explicitly include machine-verifiability of the output as part of the definition (?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?). Only one group of authors attempted a formal definition (?; ?).
@@
In this category, we reviewed 30 papers. The target formalisms include first order logic (?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?); declarative programming language, Prolog (?; ?; ?) and answer set programming (ASP) (?; ?; ?; ?). Another main line of work focuses on translating problems from natural language into temporal logics such as LTL (?) and CTL (?). These languages extend classical logic with temporal operators to reason about properties that hold over time, with LTL describing linear sequences of states and CTL describing branching time structures where multiple futures are possible from each state. There exist a significant number of recent works focused on translating natural language descriptions to these languages (?; ?; ?; ?; ?; ?; ?; ?; ?; ?); one the reasons of popularity of temporal logics is its relevance in industrial applications.
Only 8 of reviewed papers explicitly used the term autoformalization, and all of them defined it informally—typically as the translation of informal or natural language into formal representations (?; ?; ?; ?; ?; ?; ?; ?). Among the remaining works, one paper formally defines the task (?), while the rest describe it informally as translation from natural language into a specific formalism. Notably, even though these papers do not use the term autoformalization, their definitions are consistent with those that do.
@@
In the planning category, we included 16 papers. None of them use the term “autoformalization”. One of the works provides a formal definition of the task (?) as defining a mapping function generating a world model from a natural language description. The rest of the works define the task informally (?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?; ?). In some cases the task may encompass generating both domain and problem files (?), while in others only PDDL snippets are being instantiated, e.g., goals in  (?). A related but separate line of research involves using LLMs as planners themselves, generating plans from PDDL inputs. This direction falls outside the scope of the present work.
@@
Several works that we surveyed (?; ?; ?; ?) focus on translating natural language descriptions of real-world domains into ontologies using the Web Ontology Language (OWL) (?), a formal language used to define and share structured, machine-readable knowledge in terms of individuals, concepts, and relationships between them. Some other works produce ontologies expressed in RDF or RDFS (?). RDF allows us to express Knowledge Graphs, where nodes represent named entities, node labels represent their individual properties (called “types” or “classes”), and labelled edges represent relations between them. RDFS enriches RDF with information about subsumption, domain, and range relations between classes. Both RDF and RDFS can be seen as subsets of OWL. Works such as (?; ?) translate natural language to RDFS and RDF ontologies, respectively. Others express ontologies in ad-hoc languages (?; ?), which seem to be expressible in OWL or RDFS. In all these works, syntactic validity is checked automatically, and semantic is checked manually, or by running certain conjunctive queries on the produced ontology. None of the reviewed works in this category employs the term autoformalization to denote the translation task.
@@ 4 Challenges and Open Problems
While there is a growing body of work aimed at advancing autoformalization (?; ?; ?; ?; ?; ?; ?; ?; ?), several fundamental challenges still offer exciting research opportunities that will shape the field’s future development. We next provide a selective overview of these challenges and opportunities.